# A Qubit as a Bridge Between Statistical Mechanics and Quantum Dynamics

### v1

@@ VI.1 Toward Interactions
Each site $i$ hosts a spin described by the Pauli operators $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}$. In this section, we focus on an interaction Hamiltonian involving only the $\sigma^{z}$ operators:
@@
where $J/2$ sets the nearest-neighbor interaction strength. The Pauli matrices are given by
[…]
@@
This reproduces the same polynomial form as in the non-interacting case with $N-1$ such zeros lying at $y=-1$, still being outside the thermal domain.
@@
|$\ket{\Psi(0)}=\prod_{j=1}^{N}\ket{\psi_{j}(0)},\quad\ket{\psi_{j}(0)}=\frac{1}{\sqrt{2}}\left(|\uparrow_{j}\rangle+|\downarrow_{j}\rangle\right),$ (34)
@@
$\mathcal{L}(y)=\frac{1}{2^{N}}\sum_{\{s_{j}=\pm 1\}}e^{iEt/\hbar}=\left(\frac{1+y}{2}\right)^{N-1}.$ (35)
@@
This expression shares the same analytic structure as the thermal partition function and exhibits a zero at $y=-1$. Thus, under certain interactions and boundary conditions, the rate function governing the collective behavior of the system retains the same features as observed in the single-qubit case.
VII Conclusion
@@
The central idea unifying the thermal and dynamical descriptions is that both the partition function and the Loschmidt amplitude can be expressed as evaluations of a single analytic function $\mathcal{L}(y)$ along different paths in the complex $y$-plane. In thermal physics, $y=e^{\beta E}$ lies on the positive real axis, while in quantum dynamics, $y=e^{iEt/\hbar}$ traces the unit circle. This analytic continuation shows that dynamical features such as orthogonality and singularities in dynamical free energy are governed by the same underlying zero structure that also encodes thermal properties.
@@
The qubit-based approach presented here is both accessible and rigorous, and thus can easily be incorporated in a statistical mechanics or a quantum mechanics course. This framework opens a pedagogical entry point for introducing advanced concepts such as dynamical quantum phase transitions, large deviation theory, and more complex phenomena in interacting systems using only familiar tools from undergraduate physics. Such advanced topics can further be assigned as projects to the students. As the connection between equilibrium and nonequilibrium physics continues to unfold, the deceptively simple qubit serves as a reminder that deep insights often emerge from minimal models.
Acknowledgements MK thanks Axis Bank for their financial support.
@@ A.1 Mandelstam–Tamm Bound
Taking $B=H$ and using the Heisenberg equation $[A,H]=-i\hbar\frac{dA}{dt}$, we obtain
@@
Choosing $A=|\psi_{0}\rangle\langle\psi_{0}|$, the projection operator onto the system’s initial state, the return probability $P(t)=|\langle\psi_{0}|\psi(t)\rangle|^{2}$ corresponds to the expectation value of $A$ in the time-evolved state. The variance of $A$ is then given by $P(1-P)$. Substituting into Eq. (41) gives
@@
For the qubit initialized in the cat state (Eq. 3) and evolving under the Hamiltonian in Eq. 1, the energy variance is
@@ A.2 Margolus–Levitin Bound
Taking the real part and applying the inequality $\cos\theta\geq 1-\frac{2}{\pi}\theta-\frac{2}{\pi}\sin\theta$, for $0\leq\theta\leq\pi$, we obtain
@@
In our qubit model (see, Eqs. 1,  3 and 47), we have
@@
which again coincides with $\tau_{\perp}$, indicating that the Margolus–Levitin bound is also saturated in this model.

### v2

@@ VI.1 Toward Interactions
Each site $i$ hosts a spin described by the Pauli operators $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z},$ which are defined as
In this section, we focus on an interaction Hamiltonian involving only the $\sigma^{z}$ operators:
@@
where $J/2$ sets the nearest-neighbor interaction strength.
This Hamiltonian is diagonal in the basis of product eigenstates of the $\sigma_{i}^{z}$ operators. Since each spin can take eigenvalues $\pm 1$, the Hilbert space consists of $2^{N}$ distinct configurations of the form $\{s_{1},s_{2},\dots,s_{N}\}$, where $s_{i}=\pm 1$ is the eigenvalue of $\sigma_{i}^{z}$ at site $i$. The states $|\uparrow\rangle$ and $|\downarrow\rangle$ refer to the eigenstates of $\sigma^{z}$ with eigenvalues +1+1 and 1-1, respectively. A configuration is thus a sequence of $N$ such eigenstates. The corresponding energy eigenvalue of the chain is
@@
This reproduces the same polynomial form as in the non-interacting case with N1N-1 such zeros lying at y=1y=-1, still being outside the thermal domain.
@@
[…]
@@
$\mathcal{L}_{N}(y)=\frac{1}{2^{N}}\sum_{\{s_{j}=\pm 1\}}e^{-iEt/\hbar}=\left(\frac{1+y}{2}\right)^{N-1}.$ (37)
@@
This expression shares the same analytic structure as the thermal partition function and exhibits a zero at y=1y=-1. Thus, under certain interactions and boundary conditions, the rate function governing the collective behavior of the system retains the same features as observed in the single-qubit case.
The effect of interactions becomes visible in the periodic-chain case, which we suggest as a project in Appendix C. For an interacting system, if the zeros of the partition function pinch the positive real axis in the infinite-size limit, a thermal phase transition occurs [2, 10, 11]. Likewise, in the complex-$y$ plane, if the zeros in the infinite-size limit pinch or intersect the unit circle, the quantum evolution exhibits nonanalytic behavior in time. These singularities are examples of dynamical quantum phase transitions [12, 16, 17, 18, 19].
If the zeros intersect the $|y|=1$ unit circle at isolated points, the corresponding divergences in $f_{L}$ signal the orthogonalization of the states at those times. In other cases, the nature of the dynamical quantum phase transitions will be determined by the density of zeros.
VII Pedagogic values
VII.1 For students and instructors
A unified understanding becomes possible because both statistical mechanics and quantum dynamics fundamentally rely on combining contributions from all accessible states; it is through summation over states, justified by ergodicity, in the former and through coherent superposition in the latter. Once this common foundation is recognized, it becomes evident that the familiar equilibrium and time-dependent descriptions arise as different ways of exploring a single underlying complex function.
For students, this perspective offers a concrete and computable example in which the complex plane and analytic continuation become physically meaningful, linking equilibrium and dynamical quantities through different paths in the complex plane. They also see how ideas from complex analysis, statistical mechanics, and quantum dynamics come together within one accessible framework.
For teachers, it provides a compact, analytically transparent framework that can be integrated into courses to bridge statistical mechanics and quantum mechanics without heavy formalism, strengthening coherence across the curriculum.
VII.2 On analytic continuation
The observation that the Boltzmann factor in statistical mechanics and the unitary time evolution operator in quantum mechanics are related by analytic continuation is exploited in the path integral formulation of a many-body system [15]. In such a formulation, quantum dynamics is expressed in terms of trajectories in space-time, and thermal equilibrium is obtained by replacing real time with imaginary time ($t\to it$). This time-transformtion, called Wick rotation, converts a quantum problem with a Hamiltonian in $d$-dimensions into a classical statistical system living in $(d+1)$-dimensios. For example, the two-state problem would become a one-dimensional spin chain. See, e.g, Ref. [15].
Our approach takes a different path. By working directly with the Loschmidt amplitude, which plays a role analogous to the diagonal Green’s function in the many-body framework of Ref. [15], we sidestep the path-integral formalism entirely. This means that neither Wick rotation nor any additional mathematical apparatus need to be introduced. The method relies solely on concepts students typically encounter in standard courses in mathematical physics, statistical mechanics, and quantum mechanics. In this way, it offers an accessible means of understanding how equilibrium and dynamical descriptions are connected, without requiring tools beyond the undergraduate curriculum. Moreover, the same framework extends naturally to research-level topics, such as dynamical quantum phase transitions, in which the zeros of the Loschmidt echo in the complex-$y$ plane govern the behavior of the transition [12, 16, 17]. The problem of the Hamiltonian of Sec. VI.1 under periodic boundary conditions can be a project topic for advanced courses in statistical mechanics or quantum mechanics.
VIII Conclusion
@@
The central idea unifying the thermal and dynamical descriptions is that both the partition function and the Loschmidt amplitude can be expressed as evaluations of a single analytic function $\mathcal{L}(y)$ along different paths in the complex $y$-plane. With J-J and 0 as the two energy values, $y=e^{\beta J}$ lies on the positive real axis in thermal physics, while in quantum dynamics, $y=e^{iJt/\hbar}$ traces the unit circle. This analytic continuation shows that features such as orthogonality and singularities in dynamical free energy are governed by the same underlying zero structure that also encodes thermal properties. Throughout this work, the discussion is restricted to examples defined on finite-dimensional Hilbert spaces, where the spectrum is purely discrete, allowing this correspondence to be demonstrated transparently using elementary tools.
@@
The qubit-based approach presented here is both accessible and rigorous, and thus can easily be incorporated in a statistical mechanics or a quantum mechanics course. This framework opens a pedagogical entry point for introducing advanced concepts such as dynamical quantum phase transitions, and more complex phenomena in interacting systems using only familiar tools from undergraduate physics. Such advanced topics can further be assigned as projects to the students, see Appendix C. As the connection between equilibrium and nonequilibrium physics continues to unfold, the deceptively simple qubit serves as a reminder that deep insights often emerge from minimal models.
Conflict of interest The authors declare no conflict of interest.
@@ A.1 Mandelstam–Tamm Bound
Taking $B=H$ and using the Heisenberg equation $A,H]=-i\hbar\frac{dA}{dt}$, we obtain
@@
Choosing $A=|\psi_{0}\rangle\langle\psi_{0}|$, the projection operator onto the system’s initial state, the return probability $P(t)=|\langle\psi_{0}|\psi(t)\rangle|^{2}$ corresponds to the expectation value of $A$ in the time-evolved state. The variance of $A$ is then given by P(1P)P(1-P). Substituting into Eq. (43) gives
@@
For the qubit initialized in the cat state (Eq. (3)) and evolving under the Hamiltonian in Eq. (1), the energy variance is
@@ A.2 Margolus–Levitin Bound
Taking the real part and applying the inequality
$\cos\theta\geq 1-\frac{2}{\pi}\theta-\frac{2}{\pi}\sin\theta,{\textrm{ for }}0\leq\theta\leq\pi,$ (51)
we obtain
@@
In our qubit model (see, Eqs. (1),  (3) and (49)), we have
@@
which again coincides with $t_{c}$, indicating that the Margolus–Levitin bound is also saturated in this model.
A.2.1 Proof of inequality, Eq. (51)
To prove the inequality given in Eq. (51), let us consider
$k(\theta)=\frac{1-\cos\theta}{\theta+\sin\theta},$
Since
$\frac{dk}{d\theta}=\frac{\theta\sin\theta}{\theta+\sin\theta}>0,{\textrm{ for }}\theta\in[0,\pi],$
$k(\theta)$ is monotonically increasing function at least in the range $\theta\in[0,\pi],$ and
$k(\theta)=\left\{\begin{array}[]{ll}0&{\textrm{ for }}\theta=0\\ 2/\pi&{\textrm{ for }}\theta=\pi\end{array}\right$. (58)
Therefore $k(\theta)\leq 2/\pi,$ or, $\cos\theta\geq 1-\frac{2}{\pi}\theta-\frac{2}{\pi}\sin\theta.$
@@ A.3 Scaling of QSLs in Many-Body Systems
Appendix B Repeated Measurement and the Quantum Zeno Effect
The survival probability is the probability that a system remains in the state in which it was initially prepared. We now determine the survival probability over a total time $\tau$ when the system is subjected to repeated measurements of the initial state at intervals of duration $\tau/n$ [7, 8, 9].
The survival probability $P(t)$ has the short-time form
$P(t)\approx\exp(-\alpha t^{2}),$ (59)
as obtained from the early time behaviour given by Eqs. (17) and (14). This quadratic dependence reflects the general quantum-mechanical result that departures from the initial state occur only to second order in time.
According to the collapse postulate of quantum mechanics, each measurement projects the state onto the initial state whenever the system is found there, after which it again evolves for a time $\tau/n$ according to the same probability law $P(t)$ of Eq. (59). Because each short-time evolution is independent, the survival probability after $n$ such cycles is
$P_{n}(\tau)=[P(\tau/n)]^{n}=\left[e^{-\alpha\,\tau^{2}/n^{2}}\right]^{n}=e^{-\alpha\tau^{2}/n}.$ (60)
In the limit of increasingly frequent measurements,
$\lim_{n\to\infty}P_{n}(\tau)=1,$ (61)
so the system remains in its initial state with probability $1$. Thus, frequent measurements inhibit the system’s evolution away from its initial state. This phenomenon is known as the quantum Zeno effect.
Appendix C Exercises
A possible project topic:
Consider the Ising model of Section VIA with periodic boundary conditions. Explore the features of phase transitions.
This project naturally leads to the recently developed notion of dynamical quantum phase transitions (DQPT) [12, 16, 17, 18, 19]. In this example, the orthogonalization is preempted by a DQPT singularity in $f_{L}$ at a critical time $t_{c}$. There is no $\tau_{\perp}$.
The following homework problems can be assigned:
  1. 1.
    A two-level system with a degenerate excited state (two-fold degeneracy). Analyze the time evolution of the return probability, and connect its periodicity to the location of the zero.
  2. 2.
    Repeat the above analysis for a harmonic oscillator with the initial state prepared in an equal superposition of the first three energy eigenstates. Note that there are more than one zero. Generalize the result to an initial state with an arbitrary number of superposed energy levels.